OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points and having the given slope. A(5,5), M=3

OpenStudy (cwrw238):

the general equation is y-y1 = m(x-x1) where m = slope and (x1, y1) is a point on the line

OpenStudy (anonymous):

so what is the answer

OpenStudy (cwrw238):

so plug in your values and rearrange to standard form

OpenStudy (anonymous):

i dont know how

OpenStudy (cwrw238):

wellas a first step :- instead of x1 put 5, instead of y1 put 5 and instead of m put 3 what do you get?

OpenStudy (anonymous):

y-5=3x-15

OpenStudy (cwrw238):

great now can you rearrange this to standard form?

OpenStudy (anonymous):

how?

OpenStudy (cwrw238):

standard form in US may be different to Uk in Uk it would be y - 5 = 3x - 15 subtract 3x to both sides y - 5 - 3x = -15 add 5 to both sides y - 3x = -10

OpenStudy (cwrw238):

try googling it

OpenStudy (anonymous):

thanks!!!!!

OpenStudy (cwrw238):

i think the x should come before th e y so its -3x + y = -10

OpenStudy (anonymous):

ok thanks